OpenStudy (gio333):

The slope of a line is –2 and its y-intercept is (0, 3). What is the equation of the line that is parallel to the first line and passes through (2, 2)?

geerky42 (geerky42):

use this formula: \(y-y_1=m(x-x_1)\) parallel lines have same slope, so second line will have slope of -2 for point, plug in 2 for x1 and 2 for y1

geerky42 (geerky42):

is this clear?

OpenStudy (gio333):

yeah

geerky42 (geerky42):

ok, so after plug in, simplify it to y=mx+b. what did you get?

OpenStudy (gio333):

y=-2x+6

geerky42 (geerky42):

right! good job :)

OpenStudy (gio333):

thanks